your bank pay 5 percent annual interest compounded semianually on your savings account. you dont expect to add to the current balance of $2,700 over the next four year. how much money can you expect to have at the end of this period?
Answer #1

This question requires application of basic time value of money function, according to which:

FV = PV * (1 + r)n

PV = $2,700

r = 5%/2 = 2.5% (semi-annually)

n = 4 * 2 = 8 semi-annual periods

FV = $2,700 * (1 + 2.5%)8

FV = $2,700 * 1.2184

FV = $3,289.69